# Better Understandings and Configurations in MaxSAT Local Search Solvers via Anytime Performance Analysis

### v1

@@ 3.2 Anytime Performance
Figure 2(a) shows that NuWLS outperforms the other solvers for both weighted and unweighted MaxSAT. BandMax and MaxFPS show similar performance, but by examining the “anytime” analysis we can see that MaxFPS slightly outperforms BandMax regarding within a time limit of $10s$, and the order of their performances alter as the optimization time increases for the weighted MaxSAT. Furthermore, detailed analysis of anytime performance for specific instances can reveal interesting findings. Although NuWLS demonstrates advantages over other solvers considering ECDFs aggregated across the entire set of tested problem instances, it does not exhibit superiority on $15$ “decision-tree” instances after being trapped in local optima around $10s$, as shown in Figure 2(b). On the other hand, while BandMax obtains similar performance to SATLike for $13$ “ParametricRBACMaintenance” instances when cpu time is less than $100s$, it outperforms the other solvers afterward. Compared to the fixed-budget performance assessment, these anytime performance observations offer more insights into the behavior of algorithms for specific problem instances or sets of instances. These observations can be valuable for enhancing algorithm design.
Figure 4: Heatmap illustrating the ECDFs for individual instances of four benchmark tracks. It follows the layout of Figure 1, but the color depicts the ECDF values achieved by the solvers.
@@
Table 3: Aggregated ECDFs of the solvers for each benchmark track. The ECDF value for each run on an instance is normalized by the number of considered cutoff times, i.e., $100$. The presented values are average across the corresponding instances.
@@
As mentioned previously, ECDF is with a ratio scale that allows aggregation of values across multiple optimization times to obtain quantitative assessments for each instance, which is the so-called (approximated) area under the ECDF curve (AUC) Ye et al. (2022). Table 3 presents the aggregated ECDFs for the four benchmark tracks, showing that the order of the solver’s assessment based on ECDF remains the same as the one based on scores, as presented in Table 2. When analyzing the behavior in specific instances, Figure 4 shows higher variances in ECDFs for each instance compared to the ones in scores (see Figure 1). This measn that instead of having a large set of scores higher than $0.9$, as spread across the yellow area of Figure 1, we can differentiate the solvers’ performance using ECDF for most instances, except those in which no feasible solution is obtained. Furthermore, high-variance ECDFs list across different instances for each solver. Moreover, the difference in such convergence process can be quantitatively estimated using ECDFs. Due to the space limit, we are unable to provide a table with complete ECDF values, but all the data is available in our repository.
4 Hyperparameter Optimization
Contemporary algorithms, such as heuristics, evolutionary algorithms, machine learning models, etc., are usually parametric. Therefore, parameter settings are essential for the competitive performance of algorithms. Before the development of automatic tools, parameter settings were manually chosen based on experts’ experience or repetitive testing using Grid Search. Nowadays, the tools such as SMAC Lindauer et al. (2022) and irace López-Ibáñez et al. (2016) have been commonly applied for automatic HPO. While practical scenarios may propose various requirements for HPO, we address in this paper the algorithm configuration (AC) problem, which meets the requirement of tuning MaxSAT solvers for a set(s) of problem instances. The Algorithm Configuration Lindauer et al. (2022) aims at searching a well-performance configuration, i.e., parameter setting, $\lambda\in\Lambda$ of an algorithm $A$ across a set of problem instances $\{i_{1},\ldots\}\subset\mathcal{I}$:
@@ 4 Hyperparameter Optimization
where the cost function $e$ is usually set as the best-found solutions’ quality intuitively when tuning MaxSAT LS solvers. However, in this section, we introduce ECDF as an alternative candidate for $e$ and compare the results obtained by using these two different mechanisms as the cost function for tuning MaxSAT LS solvers.
@@
We use the AC arcade of SMAC3  [^1] https://github.com/automl/SMAC3 to compare the results of tuning fixed-budget performance and anytime performance, namely, the quality of best-found solutions (Best-f) and ECDFs, respectively. Using Best-f as the cost function, it straightforwardly calculates the fitness of the best-found solution obtained within a given cutoff $t_{B}$. As for calculating ECDF, we form the a set $T$ of $50$ optimization times chosen within a range of $[0.1s,t_{B}]$ following a logarithmic scale. In Section 3, we prefer solvers to obtain high-quality solutions in less time when assessing anytime performance. However, the time budget $t_{B}$ for executing each run of configurations is usually relatively small for HPO scenarios, e.g., $t_{B}=100s<300s$ in this work. Therefore, following a bias toward the configurations that have the potential to yield better solutions when provided with an additional time budget, we use $t_{B}-T$ as the set of optimization times to calculate ECDFs. Regarding the target (i.e., solution quality) set $\Phi_{i}$ for each instance $i$, we start with a set $F_{i}$ of five fitness values chosen from $[f_{\text{init\_min}},f_{\text{init\_max}}]$ based on the linear scale, where $f_{\text{init\_min}}$ and $f_{\text{init\_max}}$ are the fitness of the best and the worst solutions visited by the initial configuration. When a configuration obtains a better best-found solution for an instance during the tuning process, the corresponding best-found fitness value will be added to $F_{i}$. In addition, the ECDF of a configuration for a given instance is calculated by aggregating the values across $50$ optimization times, without normalizing it by the size of the solution fitness set. In practice, we use a modified $\text{ECDF}^{\prime}=-\text{ECDF }\cdot\mid\Phi_{i}\mid$ measuring $e$ of configurations for each instance $i$. By using $\text{ECDF}^{\prime}$, it decreases accordingly when a better configuration achieves a better best-found solution, thereby increasing the size of $\Phi$. In this way, tuning Best-f and ECDF both are both framed as minimization tasks for HPO. Though computing ECDF is more complex than computing Best-f, in practice, this computing takes linear time, and its time consumption is negligible compared to the time budget allocated for each run of a configuration.
@@
NuWLS SATLike
Instances Best-f ECDF Best-f ECDF
MSE23-w 0.386 0.450 0.139 0.237
0.398 0.455 0.184 0.246
0.409 0.456 0.325 0.329
0.421 0.465 0.328 0.410
0.457 0.503 0.329 0.412
Avg. 0.414 0.466 0.261 0.327
MSE23-uw 0.421 0.432 0.307 0.372
0.450 0.473 0.337 0.476
0.457 0.505 0.345 0.483
0.483 0.531 0.409 0.510
0.495 0.591 0.456 0.549
Avg. 0.461 0.506 0.371 0.478
[…]
(a) NuWLS
(b) SATLike
@@
Figure 5: ECDFs of the NuWLS and SATLike configurations obtained by tuning for Best-f and ECDF for two sets of instances: Left: MSE23-w and Right: MSE23-uw. Results plotted in one line are from five configurations obtained by independent runs of SMAC.
NuWLS SATLike
Instances Best-f ECDF Best-f ECDF
MSE23-w 0.833 0.855 0.495 0.676
0.835 0.858 0.515 0.684
0.842 0.862 0.807 0.834
0.848 0.870 0.816 0.903
0.859 0.883 0.835 0.907
Avg. 0.843 0.866 0.694 0.801
MSE23-uw 0.814 0.823 0.566 0.654
0.818 0.837 0.577 0.830
0.829 0.845 0.584 0.837
0.840 0.856 0.733 0.844
0.846 0.891 0.757 0.860
Avg. 0.829 0.850 0.643 0.805
Table 5: Aggregated scores and the average of five NuWLS configurations and five SATLike configurations.
Anytime Performance Leads to Better Configurations.
We first compare the anytime performance of the obtained configurations. Table 4 presents the aggregated ECDFs across weighted and unweighted MaxSAT instances, and Figure 5 demonstrates the mean and $95\%$ confidence interval of ECDFs of five configurations along cpu time. Since we focus on the impact of the cost function, we analyze the comparison results separately for NuWLS and SATLike. These results are computed based on the corresponding ten configurations for each benchmark track. According to Table 4 and Figure 5, for both weighted and unweighted tracks, the optimal configuration among the ten ones is obtained by using ECDF as the cost function, and the results obtained by tuning for ECDF generally outperform the ones of tuning for Best-f. More interestingly, as observed in Table 5, tuning for ECDFs enables obtaining configurations that can achieve better scores than those obtained by tuning for the fitness of the best-found solutions. This suggests that using ECDF as the cost function can lead to better configurations in terms of both anytime performance and best-found solution quality.
One reason for this finding is that ECDFs can provide dense search space for HPO compared to Best-f, as shown in Figure 1 and Figure 4. For example, within a given time budget $t_{B}$, even though two different configurations $\lambda_{1}$ and $\lambda_{2}$ may obtain the same best-found solution for an instance $i$ using different time $t\leq t_{B}$, Best-f will deliver identical information to HPO tools, resulting in a sparse search space, but ECDF can distinguish these two configurations.
@@
In addition, when conducting HPO, the time budget $t_{B}$ is usually set as a relatively small value due to time limits. For example, in our tuning process of SMAC, we allocated a budget of $100s$ for each configuration. However, when assessing the final obtained configurations, the budget is set as $300s$. In such scenarios where configurations are allocated with limited time budgets, via considering multiple cutoff times, ECDF can robustly estimate the potential ability of a configuration to achieve better solutions within a larger time budget. On the other hand, using Best-f as the cost function may be misled by fortunate results obtained at a specific cutoff time.
5 Conclusion
@@
We have introduced ECDF for anytime performance analysis of MaxSAT local search solvers. Our assessments have been conducted for the four state-of-the-art solvers, namely BandMax, MaxFPS, NuWLS, and SATLike. We illustrate that ECDF can measure solvers’ anytime performance regarding multiple cutoff times, provide quantitative assessments with a ratio scale that can be aggregated across multiple instances, and differentiate the solvers that are considered identical in terms of fixed-budget performance. Although we have presented the application in the context of MaxSAT local search, we expect that the applied techniques can be transparent to the hybrid methods integrating complete and incomplete solvers. Our experiments show that ECDF can serve as a valuable metric for hyperparameter optimization to obtain better configurations of local search solvers. Additionally, by utilizing customized sets of time budgets and solution fitness to calculate ECDF, we anticipate that ECDF can serve as a valuable tool for researchers to comprehend the behavior of algorithms and improve their design for specific scenarios. Furthermore, we will study the algorithm portfolio configuration, which has a significant impact on MaxSAT, considering anytime performance of solvers.
6 Data Availability
The source code and raw data of the tested solvers, code and reproducible output of SMAC scenarios, post-processing results presented in the paper, the performance of tuning BandMax, and example data considering Function Evaluations as the time metric are publicly available in our repository: https://github.com/AcademicSubmission/IJCAI2024-p3562/.

### v2

@@ Introduction
Figure 3(a) shows that NuWLS outperforms the other solvers for both weighted and unweighted MaxSAT. BandMax and MaxFPS show similar performance, but by examining the “anytime” analysis, we can see that MaxFPS slightly outperforms BandMax within a time limit of $10s$, and the order of their performances alter as the optimization time increases for the weighted MaxSAT. Furthermore, detailed analysis of anytime performance for specific instances can reveal exciting findings. Although NuWLS demonstrates advantages over other solvers considering ECDFs aggregated across the entire set of tested problem instances, it does not exhibit superiority on “decision-tree” instances after being trapped in local optima around $10s$, as shown in Figure 3(b). On the other hand, while BandMax obtains similar performance to SATLike for “ParametricRBACMaintenance” instances when cpu time is less than $100s$, it outperforms the other solvers afterward. Compared to the fixed-budget performance assessment, these anytime performance observations offer more insights into the behavior of algorithms for specific problem instances and can be valuable for enhancing algorithm design.
Figure 4: Heatmap illustrating aggregated ECDFs for individual instances. It follows the layout of Figure 1, but the color depicts the ECDF values achieved by the solvers. Detailed results regarding groups of instances are available in Appendix B.
@@
Table 3: Aggregated ECDFs for each benchmark track. The ECDF value for each run on an instance is normalized by the number of considered cutoff times, i.e., $100$. The presented values are average across the corresponding instances.
@@
As mentioned previously, ECDF has a ratio scale that allows aggregation of values across multiple optimization times to obtain quantitative assessments for each instance, which is the so-called (approximated) area under the ECDF curve (AUC) (Ye et al. 2022). Table 3 presents the aggregated ECDFs for the four benchmark tracks, showing that the order of the solver’s assessment based on ECDF remains the same as the one based on scores, as presented in Table 2. When analyzing the behavior in specific instances, Figure 4 shows higher variances in ECDFs for each instance compared to the ones in scores (see Figure 1). Specifically, instead of having a large set of scores higher than $0.9$ in Figure 1, ECDFs are in a wide variation of values as shown in Figure 4. We also provide the average standard deviations of scores and ECDFs across all tested instances in Appendix A, which shows that the standard deviations obtained from ECDFs are much greater than the scores. Moreover, for the instances where solvers can obtain the same best-found solutions, ECDFs can distinguish solvers’ performance by quantitatively estimating the convergence process. For example, Figure 4 can show that NuWLS converge faster for the instances plotted in yellow.
Hyperparameter Optimization
Contemporary algorithms, such as heuristics, evolutionary algorithms, machine learning models, etc., are usually parametric. Therefore, parameter settings are essential for the competitive performance of algorithms. Before the development of automatic tools, parameter settings were manually chosen based on experts’ experience or repetitive testing using Grid Search. Nowadays, the tools such as SMAC (Lindauer et al. 2022) and irace (López-Ibáñez et al. 2016) have been commonly applied for automatic HPO. While practical scenarios may propose various requirements for HPO, we address in this paper the algorithm configuration (AC) problem, which meets the requirement of tuning MaxSAT solvers for a set(s) of problem instances. AC (Lindauer et al. 2022) aims at searching a well-performance configuration, i.e., parameter setting, $\lambda\in\Lambda$ of an algorithm $A$ across a set of problem instances $\{i_{1},\ldots\}\subset\mathcal{I}$:
@@
where the cost function $e$ is usually set as the best-found solutions’ quality intuitively when tuning MaxSAT SLS solvers. However, in this section, we introduce AUC as an alternative candidate for $e$ and compare the results obtained by using these two different mechanisms as the cost function for tuning MaxSAT SLS solvers. For ease of reading, we denote AUC and aggregated ECDFs as the metrics used in HPO and experimental comparisons, respectively.
@@
We use the AC arcade of SMAC3  [^2] https://github.com/automl/SMAC3 to compare the results of tuning fixed-budget performance and anytime performance. Best-f, which calculates the fitness of the best-found solution obtained within a given cutoff $t_{B}$, is a straightforward option for the cost function of tuning fixed-budget performance. In addition, we apply $-score$ in Eq. 1, which is denoted as Norm-f, as another cost function of tuning fixed-budget performance. To calculate scores, we use the best-found solutions obtained in the Performance Assessment section. Note that this function can be adopted for competitions but is not doable for the practical scenarios lacking baselines, i.e., the known best-found solutions. As for calculating the cost function of anytime performance approximated AUC, we form a set $T$ of $50$ optimization times chosen within a range of $[0.1s,t_{B}]$ following a logarithmic scale. Since the time budget $t_{B}$ for executing each run of configurations is usually relatively small for HPO scenarios, e.g., $t_{B}=100s<300s$ in this work, following a bias toward the configurations that have the potential to yield better solutions when provided with an additional time budget, we use $t_{B}-T$ as the set of optimization times to calculate AUCs. Regarding the target (i.e., solution quality) set $\Phi_{i}$ for each instance $i$, we start with a set $F_{i}$ of five fitness values chosen from $[f_{\text{init\_min}},f_{\text{init\_max}}]$ based on the linear scale, where $f_{\text{init\_min}}$ and $f_{\text{init\_max}}$ are the fitness of the best and the worst solutions visited by the initial configuration. When a configuration obtains a better best-found solution for an instance during the tuning process, the corresponding best-found fitness value will be added to $F_{i}$. In addition, the AUC of a configuration for a given instance is calculated by aggregating the values across $50$ optimization times. In practice, we use a modified $\text{AUC}^{\prime}=-\text{AUC }\cdot\mid\Phi_{i}\mid$ measuring $e$ of configurations for each instance $i$. $\text{AUC}^{\prime}$ decreases accordingly when a better configuration achieves a better best-found solution, thereby increasing the size of $\Phi$. In this way, all of the tuning scenarios are framed as minimization tasks for HPO. Though computing AUC is more complex than computing Best-f and Norm-f, in practice, this computing takes linear time, and its time consumption is negligible compared to the time budget allocated for each run of a configuration.
We have tuned the four solvers for the MSE23-w and MSE23-uw benchmark tracks by using SMAC with three different cost functions. We conduct five runs of each SMAC setting with the cpu time budget of $60,000s$, following the suggestion in the work of NuWLS (Chu, Cai, and Luo 2023). Twenty percent of the instances are randomly selected for training. The obtained configurations of solvers are compared based on scores and AUCs across all the instances. Note that the computation of scores and ACUs follows the same settings as the Performance Assessment section.
BandMax MaxFPS NuWLS SATLike
Best-f Norm-f ECDF Best-f Norm-f ECDF Best-f Norm-f ECDF Best-f Norm-f ECDF
Results on MSE23-w
Score 0.857 0.850 0.858 0.844 0.850 0.848 0.795 0.809 0.818 0.664 0.564 0.758
ECDF 0.423 0.405 0.423 0.424 0.423 0.430 0.431 0.450 0.479 0.288 0.225 0.353
Results on MSE23-uw
Score 0.815 0.799 0.814 0.776 0.773 0.769 0.824 0.827 0.846 0.617 0.717 0.790
ECDF 0.549 0.532 0.548 0.505 0.503 0.496 0.488 0.491 0.528 0.373 0.423 0.477
Table 4: Average scores and ECDFs of five configurations obtained by each setting. Larger values indicate better performances.
Anytime Performance can Lead to Better Configurations.
@@
We first compare the anytime performance of the obtained configurations for eight scenarios (4 solvers $\times$ 2 benchmark tracks). According to Table 4, using AUC as the cost function can obtain the best anytime performance for six out of eight scenarios, except for BandMax and MaxFPS in MSE23-uw. We can observe in Figure 5a-b that the anytime performance of the tuned BandMax and MaxFPS configurations are close to each other, and the results of using AUC as the cost function generally obtain slight advantages along CPU time for BandMax. When comparing the fixed-budget performance, i.e., scores, using AUC as the cost function obtains the best results for five out of eight scenarios in Table 4. Note that Wilcoxon signed-rank tests indicate that the results of using AUC as the cost function significantly differ from the others. P-values are provided in Appendix C.
Overall, our results suggest that tuning anytime performance can lead to better configurations in terms of both anytime performance and best-found solution quality. Specifically, when tuning NuWLS and SATLike, using AUC as the cost function obtains significant advantages ($4\%$ and $15\%$ by average) compared to the second-best results. For the five out $16$ scenarios (in Table 4) that tuning AUC does not obtain the best results, the obtained results are only $0.6\%$ worse than the best ones by average.
(a) BandMax
(b) MaxFPS
(c) NuWLS
(d) SATLike
@@
Figure 5: Aggregated ECDFs of the configurations obtained by tuning for Best-f, Norm-f, and ECDF for Left: MSE23-w and Right: MSE23-uw. Results plotted in one line are from five configurations obtained by independent runs of SMAC.
Reasons for the Better Results.
@@
One reason for the findings is that AUCs can provide dense search space for HPO. For example, within a given time budget $t_{B}$, when two configurations $\lambda_{1}$ and $\lambda_{2}$ obtain the same best-found solution for an instance $i$ using different time $t\leq t_{B}$, Best-f and Norm-f will deliver identical information to HPO tools, but AUC can distinguish these two configurations. Another reason is that AUCs are normalized values ranging between 00 and $1$ compared to Best-f. Although Norm-f is also normalized, its values range in a smaller domain for each instance. In contrast, AUCs have a wider variation of values. Note that we have also tested using the average of AUCs and Norm-f values across 20 instances as the cost function each time when SMAC evaluates a configuration. However, according to the results in Appendix D, normalization across multiple instances does not help configure MaxSAT solvers. In addition, when conducting HPO, the time budget $t_{B}$ is usually set as a relatively small value due to time limits. For example, in our tuning process of SMAC, we allocated a budget of $100s$ for each configuration. However, when validating the obtained configurations, the budget is $300s$. In the scenarios where configurations are allocated with limited time budgets, by considering multiple cutoff times, AUCs can robustly estimate the potential ability of a configuration to achieve better solutions. On the other hand, using Best-f or Norm-f as the cost function may be misled by fortunate results obtained at a specific cutoff time.
Conclusion
@@
We have introduced ECDF for anytime performance analysis of MaxSAT SLS solvers. Our assessments have been conducted for the four state-of-the-art solvers, namely BandMax, MaxFPS, NuWLS, and SATLike. We illustrate that ECDF can measure solvers’ anytime performance regarding multiple cutoff times, provide quantitative assessments with a ratio scale that can be aggregated across multiple instances, and differentiate the solvers that are considered similar in terms of fixed-budget performance. Our experiments show that the AUC can serve as a competitive metric for hyperparameter optimization. Compared to the traditional fixed-budget cost functions, the AUC can help achieve configurations that are, on average, around $10\%$ better than the second-best ones in 11 out of 16 tested scenarios. In the remaining scenarios, its results are, on average, only $0.6\%$ worse than the best ones. Although we have presented the application in the context of MaxSAT SLS, we expect that the applied techniques can be transparent to the hybrid methods integrating complete and incomplete solvers. Furthermore, we have observed that the choice of cost functions can affect the results of HPO. Therefore, we plan to study multi-objective HPO and explore the effect of cost functions. Also, we will study the algorithm portfolio configuration by considering the solvers’ anytime performance.
Acknowledgements
This work is supported by the National Key R&D Program of China (2023YFA1009500), the National Natural Science Foundation of China (62202025), the Beijing Natural Science Foundation (L241050), Young Elite Scientist Sponsorship Program by CAST (YESS20230566), CCF-Huawei Populus Grove Fund (CCF-HuaweiFM2024005), and Fundamental Research Fund Project of Beihang University.